i am homeschooled i do not have a text book curriculum. With a certain tail wind a jet aircraft arrives at its destination, 1,890 miles away, in three hours. Flying against the same wind, the plane makes the return trip in 3 3/8 hours. find the wind speed and the planes airspeed.
Answer by stanbon(75887)

With a certain tail wind a jet aircraft arrives at its destination, 1,890 miles away, in three hours. Flying against the same wind, the plane makes the return trip in 3 3/8 hours. find the wind speed and the planes airspeed.
With the wind Data:
distance = 1890 miles; time=3 hrs; rate=1890/3=630mph
Against the wind Data:
distance = 1890 miles: time= 3.375 hrs; rate=1890/3.375=560mph
airspeed+wind=630mph
airspeed-wind=560mph
Add the two equations to get:
2(airspeed)=1190mph
Airspeed=595mph
Then wind=630-595=35mph
Cheers,
stan H.
PS: To get much further in math you need instruction.